Saddle Inc. has two types of handbags: standard and custom. The controller has decided to use a plantwide overhead rate based on direct labor costs. The president has heard of activity-based costing and wants to see how the results would differ if this system were used. Two activity cost pools were developed: machining and machine setup. Presented below is information related to the company’s operations.

Standard

Custom

Direct labor costs $52,000 $120,000
Machine hours 1,170 1,100
Setup hours 99 420


Total estimated overhead costs are $301,000. Overhead cost allocated to the machining activity cost pool is $197,000, and $104,000 is allocated to the machine setup activity cost pool.

Incorrect answer iconYour answer is incorrect.

Compute the overhead rate using the traditional (plantwide) approach. (Round answer to 2 decimal places, e.g. 12.25.)

Predetermined overhead rate

enter the predetermined overhead rate in percentages of direct labor cost rounded to 2 decimal places

% of direct labor cost

eTextbook and Media

  

  

Partially correct answer iconYour answer is partially correct.

Compute the overhead rates using the activity-based costing approach. (Round answers to 2 decimal places, e.g. 12.25.)

Machining

$enter a dollar amount per machine hour rounded to 2 decimal places

per machine hour
Machine setup

$enter a dollar amount per setup hour rounded to 2 decimal places

per setup hour

eTextbook and Media

  

  

Partially correct answer iconYour answer is partially correct.

Determine the difference in allocation between the two approaches. (Round answers to 0 decimal places, e.g. 1,225.)

Traditional costing
Standard

$enter a dollar amount rounded to 0 decimal places

Custom

$enter a dollar amount rounded to 0 decimal places

Activity-based costing
Standard

$enter a dollar amount rounded to 0 decimal places

Custom

$enter a dollar amount rounded to 0 decimal places

eTextbook and Media
Answer #1
Answer

PREDETERMINE OVERHEAD RATE = Overhead cost*100/direct labour cost

301000*100/172000
Predetermine overhead rate = 175%
COMPUTE OVERHEAD RATE UNDER ABC:
Machining = 197000/2270 = 86.78 per machine hour
set up = 104000/519 = 200.39 per set up
DETERMINE THE DIFFERENCE IN ALLOCATION BETWEEN THE TWO APPROACHES :
TRADITION COSTING
Standard (52000*175%) $ 91,000
custom (120000*175%) $210,000
ACTIVITY BASED COSTING
Standard
Machining (1170*86.78) $101,533
Setup (99*200.39) $ 19,839 $121,371
Custom
Machining (1100*86.78) $ 95,458
Setup (420*200.39) $ 84,164 $179,622